OpenStudy (anonymous):

Which pairs are solutions to the equation? 3xy + 9 = 33 A. (1, 11) and (2, 6) B. (1, 8) and (1, 11) C. (8, 1) and (2, 6) D. (1, 8) and (8, 1)

563blackghost (563blackghost):

Well 3(1)11+9=42, 3(2)6+9=45, 3(8)1+9=33 So which answer is it?

563blackghost (563blackghost):

@curtcurt14